My mother was 28 years old when I was born. The sum of our ages is 84, How old am I?
Answer by checkley71(8403)

Mother=28+x
You are x
28+x+x=84
2x=84-28
2x=56
x=28 Your age now.
28+28=56 Mom's age now.
proof
28+2*28=84
28+56=84
84=84